OpenStudy (anonymous):

what is the ordered pair for a slope y=1/4x-1? (4,?) (0,?)

OpenStudy (anonymous):

You can plug in the first coordinate there for the x to get the ? coordinate, the y coordinate. So y = (1/4) * 4-1 = 1 - 1 = 0 and y = (1/4)*4 - 1 = 0 - 1 = -1 And then you just put them back in your ordered pairs.